ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 209, 2015

Adams, T. M. and C. E. Silva On infinite transformations with maximal control of ergodic two-fold product powers

929

Artstein-Avidan, S. and B. A. Slomka On weighted covering numbers and the Levi–Hadwiger conjecture

125

Banakh, T., M. Morayne, R. Rałowski and S. Żeberski Topologically invariant σ-ideals on the Hilbert cube

715

Böhm, G. Yetter–Drinfeld modules over weak multiplier bialgebras  85
Braun, A. Noetherian completion and PI rings  487

Bruin, H., R. Fokkink and C. Kraaikamp The convergence of the generalised Selmer algorithm

803

Buckley, J., X. Massaneda and B. Pridhnani Gaussian analytic functions in the unit ball

855

Costa, A. E. and M. Farber Geometry and topology of random 2-complexes

883

Dinur, I. and V. Guruswami PCPs via the low-degree long code and hardness for constrained hypergraph coloring

611

Elekes, M. and Z. Vidnyánszky Haar null sets without G_{δ} hulls  199

Farber, M. See A. E. Costa Favre, C. and T. Gauthier Distribution of postcritically finite poly-

nomials

235

Fokkink, R. See H. Bruin Gauthier, T. See C. Favre Gavrilovich, M. and A. Hasson Exercices de style: A homotopy

theory for set theory

15

Ghasemi, S. Isomorphisms of quotients of FDD-algebras  825

Glasby, S. P., C. E. Praeger and B. Xia Decomposing modular tensor products: ‘Jordan partitions’, their parts and p-parts

215

Guruswami, V. See I. Dinur Hasson, A. See M. Gavrilovich Hong, J. and A. Kontorovich Almost prime coordinates for

anisotropic and thin pythagorean orbits

397

Huang, H., G. Liu and Y. Ye Graded elementary quasi-Hopf algebras of tame representation type 157 Iommi, G., T. Jordan and M. Todd Recurrence and transience for

suspension flows

547

Jing, Y. Semigroupoid C*-algebras and ultragraph C*-algebras  593

Jordan, T. See G. Iommi

Kleinbock, D. and K. Merrill Rational approximation on spheres  293

Kontorovich, A. See J. Hong Kozma, G. and V. Sidoravicius Lower bound for the escape prob-

ability in the Lorentz Mirror Model on ℤ^{2}

683

Kraaikamp, C. See H. Bruin Kuribayashi, K., L. Menichi and T. Naito Derived string topol-

ogy and the Eilenberg–Moore spectral sequence

745

Lindenstrauss, E. and N. d. Saxcé Hausdorff dimension and subgroups of SU(2)

335

Liu, G. See H. Huang Liu, L. Homological smoothness and deformations of generalized Weyl

algebras

949

Lücke, P. and P. Schlicht Continuous images of closed sets in generalized Baire spaces

421

J. Maříková and M. Shiota Measuring definable sets in o-minimal fields

687

Massaneda, X. See J. Buckley Menichi, L. See K. Kuribayashi Merrill, K. See D. Kleinbock Morayne, M. See T. Banakh Naito, T. See K. Kuribayashi Piatek, B. The fixed point property and unbounded sets in spaces of

negative curvature

323

Plebanek, G. On isomorphisms of Banach spaces of continuous functions

1

Pol, E. and R. Pol Note on isometric universality and dimension  187

Pol, R. See E. Pol Praeger, C. E. See S. P. Glasby Pridhnani, B. See J. Buckley Rałanski, R. See T. Banakh Richter, B. On the homology and homotopy of commutative shuffle

algebras

651

Roche-Newton, O. and M. Rudnev On the Minkowski distances and products of sum sets

507

Rudnev, M. See O. Roche-Newton Saxcé, N. d. See E. Lindenstrauss Schlicht, P. See P. Lücke Sevost’yanov, E. The Miniowitz and Vuorinen theorems for the

mappings with non-bounded characteristics

527

Shami, Z. A dichotomy for D-rank 1 types in simple theories  993

Shiota, M. See J. Maříková Sidoravicius, V. See G. Kozma Silva, C. E. See T. M. Adams

Sinapova, D. and S. Unger Scales at ℵ_{ω}  463

Slomka, B. A. See S. Artstein-Avidan Todd, M. See G. Iommi Unger, S. See D. Sinapova Vidnyánszky, Z. See M. Elekes Xia, B. See S. P. Glasby

Yadav, M. K. Class-preserving automorphisms of finite p-groups II  355

Ye, Y. See H. Huang Żeberski, S. See T. Banakh